OpenStudy (anonymous):

Find the solution of the exponential equation, correct to four decimal places. 45(1+10^(4x)=14

OpenStudy (anonymous):

45+45(10^4x)=14 450^4x=-31 1^4x = -31/450 log(1^4x)=log(-31/450) 4x=log(-31/450) x=1/4(log(-31/450)) .. i think that's right.. it's been a while

OpenStudy (anonymous):

oh wait. perhaps the first step would have been divide both sides by 45 then subtract 1 from both sides then take the log of both sides leaving you with 4x = log(14/15-1) x=1/4log(14/15-1)

OpenStudy (anonymous):

I calculated teh answer in a more efficient way lol, I graphed both sides of the equation and found the intersection, rather than doing the whole arithmetic steps